OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the point (1, 1).

OpenStudy (anonymous):

\[y=\ln(xe ^{x ^{5}})\]

OpenStudy (saifoo.khan):

satelite? i want the answer in compunt form.. with infinty and all.

OpenStudy (anonymous):

i would first rewrite this as \[\ln(x)+\ln(e^{x^5})=\ln(x)+x^5\] and go form there

OpenStudy (anonymous):

@saifoo i guess you should have \[(-\frac{1}{2},\infty)\]

OpenStudy (saifoo.khan):

and?

OpenStudy (saifoo.khan):

wht about -5/4?

OpenStudy (anonymous):

that is it. compound means "both" as far as i understand it.

OpenStudy (saifoo.khan):

can u please come back to ques..??

OpenStudy (anonymous):

sure link to it

OpenStudy (anonymous):

y=ln(xe^x^5) y=lnx+ln e^x^5 y=lnx+x^5 ln e y=lnx+x^5 dy/dx = i/x +5x^4 at point(1,1) slope m=1/1 + 5(1)^4=6 y-y1=m(x-xi) y-1=6(x-1) y=6x-6+1 y+6x-5 ans......eq of the tangent line at pt(1,1)

OpenStudy (anonymous):

did you get this kehara? any question?

OpenStudy (anonymous):

oops sorry it is y=6x-5 ans...lol